OpenStudy (caerus):

find a set of solutions,unless the statement of the exercise stipulates otherwise.

OpenStudy (caerus):

\[y^2dx-x(2x+3y)dy=0\]

OpenStudy (caerus):

@jiteshmeghwal9

OpenStudy (caerus):

i try to start this but i cant do it lol

OpenStudy (jiteshmeghwal9):

It is a homogeneous differential equation \[\frac{dy}{dx}=\frac{y^2}{2x^2+3xy}......(1)\]substitute y=vx then\[\frac{dy}{dx}=v+x \frac{dv}{dx}.......(2)\]substitute (2) in (1).

OpenStudy (jiteshmeghwal9):

\[v+x \frac{dv}{dx}=\frac{v^2}{2+3v}\]\[x \frac{dv}{dx}=\frac{-4v^2-2v}{2+3v}\]\[\frac{(2+3v)}{-4v^2-2v}dv=\frac{dx}{x}\]integrate both sides

OpenStudy (jiteshmeghwal9):

can u solve it further ?

OpenStudy (caerus):

how can i determine the problem if homogeneous or linear?

OpenStudy (jiteshmeghwal9):

If the function satisfies the condition \[f(\lambda x, \lambda y)=\lambda ^n f(x,y) \] then the the function is homogeneous

OpenStudy (caerus):

lol i didnt see, its already the miscellaneous exercises i was solving lol

OpenStudy (jiteshmeghwal9):

LOL

OpenStudy (caerus):

its the mixture of several topic problems, but im checking for linear

OpenStudy (jiteshmeghwal9):

well this one is homogeneous

OpenStudy (caerus):

(3xy-4y-1)dx +x(x-2)dy=0 <--- this one is linear right?

OpenStudy (caerus):

@jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

yes

OpenStudy (caerus):

ill open another post for your medal ^.^ we work for this..

OpenStudy (jiteshmeghwal9):

LOL